OpenStudy (anonymous):

Is the equation an example of growth or decay? y = 2.1(0.15)^x

OpenStudy (mathstudent55):

Start with x = 1. When you raise 0.15 to a larger and larger exponent, does it increase or decrease.

OpenStudy (anonymous):

Increase (:

OpenStudy (mathstudent55):

0.15^1 = 0.15 0.15^2 = 0.0225 0.15^3 = 0.003375 0.15^4 = 0.00050625 etc. As the exponent of 0.15^x increases, the 0.15^x decreases.